OpenStudy (anonymous):

Use the Zero Product Property to find the x- and y-intercepts for the following equation: y=3x^2-4x-4

OpenStudy (anonymous):

*make y=0 3x^2-4x-4=0 (3x+2)(x-2)=0 x=-2/3 & 2 (-2/3,0)&(2,0) *make x=0 3(0)^2-4(0)-4=y y=-4 (0,-4)

OpenStudy (campbell_st):

wow... and I thought Open Study frowned on giving answers

OpenStudy (anonymous):

Thanks for the help

OpenStudy (skullpatrol):

Open study does frown upon giving direct answers, but that doesn't mean people will ever stop doing it :-)